OpenStudy (anonymous):

Please as soon as possible The very rigorous medical program at a local university has a 20% drop out rate for each year. If the school admits 1,000 freshmen, how many diplomas will need to be ordered 4 years later? 656 diplomas must be ordered 410 diplomas must be ordered 316 diplomas must be ordered 240 diplomas must be ordered

OpenStudy (anonymous):

only the answer please

OpenStudy (anonymous):

correct answer gets medal

OpenStudy (anonymous):

geometric zequence in wich the common factor is 1/5

OpenStudy (kropot72):

At the end of each year the number of students is 0.8 * the initial number at the start of the year. At the end of the fourth year there will be \[1000\times 0.8\times 0.8\times 0.8\times 0.8=?\ students\]

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

@kropot72 where are u

OpenStudy (kropot72):

@pvs285 Do you have access to a calculator?

OpenStudy (anonymous):

got the answer

OpenStudy (anonymous):

410